OpenStudy (anonymous):

Solve the following equation: 3^{x}*5^{x-2} = 3^{4x} I would like a hint, not a solution.

OpenStudy (anonymous):

1) divide both sides by \[3^x\] 2) take the log of both sides to get the exponents on the ground floor instead of in the sky 3) do a bunch of algebra

OpenStudy (anonymous):

How about this strategy: Take the log base 3 of both sides

OpenStudy (anonymous):

log[3]{3^x * 5^{x-2}} = log[3]{3^{4x}}

OpenStudy (anonymous):

Using the product rule for Logs, as well as the power rule, we get:

OpenStudy (anonymous):

log[3]{3^x} + log[3]{5^{x-2}} = 4x*log[3]{3}

OpenStudy (anonymous):

x + (x-2)*log[3]{5} = 4x

OpenStudy (anonymous):

x + x*log[3]{5} - 2log[3]{5} = 4x

OpenStudy (anonymous):

Ummm

OpenStudy (anonymous):

Is this wrong?

OpenStudy (anonymous):

I want to know if the method I took above would eventually work out.

OpenStudy (anonymous):

start with \[5^{x-2}=3^{3x}\]

OpenStudy (anonymous):

Using the approach I took, I get x = ( 2*log[3]{5} / log[3]{5} - 3 )

OpenStudy (anonymous):

then take the log get \[(x-2)\ln(5)=3x\ln(3)\]

OpenStudy (anonymous):

WOW!

OpenStudy (anonymous):

then multiply out to get \[\ln(5)x-2\ln(5)=3\ln(3)x\]

OpenStudy (anonymous):

I just saw that now; the quotient rule for exponents used on 3^4x and 3^x

OpenStudy (anonymous):

OHHH That makes it so much simpler

OpenStudy (anonymous):

so \ln(5)x-3\ln(3)x=2\ln(5)\] \[(\ln(5)-3\ln(3))x=2\ln(5)\] and last step is to divide.

OpenStudy (anonymous):

oops \[\ln(5)x-3\ln(3)x=2\ln(5)\]

OpenStudy (anonymous):

then \[(\ln(5)-3\ln(3))x=2\ln(5)\]

OpenStudy (anonymous):

How do you type like that?

OpenStudy (anonymous):

and finally \[x=\frac{2\ln(5)}{ln(5)-3\ln(3)}\]

OpenStudy (anonymous):

latex. i am typing latex commands

OpenStudy (anonymous):

But aren't I using that too? I am using it in this normal box.

OpenStudy (anonymous):

actually that is the reason i started answering to practice my latex. now i am hooked. i am writing them by hand. like rolling your own

OpenStudy (anonymous):

start with \[

OpenStudy (anonymous):

end with \]

OpenStudy (anonymous):

\[test \]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

and put your equations in between try \frac{a}{b}